# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

### v3

@@ Appendix A Equivalence of Stability and List-Replicability
$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}\right]\geq 1-\frac{\delta}{2}.$ (12)
A direct consequence of Claim A.2 is that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{tn_{0}+n_{1}}}\left[A,B\right]\geq 1-\delta.$
Condition on events $A$ and $B$, and let $h^{*}$ be a stable hypothesis for $\bm{\mathcal{A}}$, as described in Eq. 10. We will show that $h^{*}$ is a candidate for output, so $\bm{\mathcal{A}}^{\prime}$ will not output “failure”. To check the first condition for output, we combine $B$ and Eq. 10 to show that
@@
$\mathcal{L}_{\bm{Q}}(h^{*})\leq\mathcal{L}_{\mathcal{D}}(h^{*})+\frac{\epsilon}{3}\leq\frac{2\epsilon}{3}.$
@@
Furthermore, $h_{o}$ also satisfies the condition $\mathcal{L}_{\bm{Q}}(h_{o})\leq\frac{2\epsilon}{3}$, so because of $A$,
$\mathcal{L}_{\mathcal{D}}(h_{o})\leq\mathcal{L}_{\bm{Q}}(h_{o})+\frac{\epsilon}{3}\leq\epsilon.$

### v4

@@ Appendix B Equivalence of global stability and list replicability
$\Pr_{{\bm{Q}}\sim\mu^{n_{1}}}\left[\left|\operatorname{loss}_{\mu}(h)-\operatorname{loss}_{\bm{Q}}(h)\right|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}\right]\geq 1-\frac{\delta}{2}.$ (13)
A direct consequence of Claim B.2 is that
$\Pr_{\mathbf{S}\sim\mu^{tn_{0}+n_{1}}}\left[A,B\right]\geq 1-\delta.$
Condition on events $A$ and $B$, and let $h^{*}$ be a stable hypothesis for $\bm{\mathcal{A}}$, as described in (11). We will show that $h^{*}$ is a candidate for output, so $\bm{\mathcal{A}}^{\prime}$ will not output “failure”. To check the first condition for output, we combine $B$ and (11) to show that
@@
$\operatorname{loss}_{\bm{Q}}(h^{*})\leq\operatorname{loss}_{\mu}(h^{*})+\frac{\epsilon}{3}\leq\frac{2\epsilon}{3}.$
@@
Furthermore, $h_{o}$ also satisfies the condition $\operatorname{loss}_{\bm{Q}}(h_{o})\leq\frac{2\epsilon}{3}$, so because of $A$,
$\operatorname{loss}_{\mu}(h_{o})\leq\operatorname{loss}_{\bm{Q}}(h_{o})+\frac{\epsilon}{3}\leq\epsilon.$